OpenStudy (anonymous):

If the graph of y=x^3+ax^2+bx-4 has a point of inflection at (1, -6), find the value of b.

jimthompson5910 (jim_thompson5910):

hint: f(x) has a point of inflection when f '' (x) = 0

OpenStudy (anonymous):

I know that y"=6x+2a but what's next?

jimthompson5910 (jim_thompson5910):

replace y'' with 0, since f '' (x) = 0

jimthompson5910 (jim_thompson5910):

replace x with the x coordinate of the point of inflection

OpenStudy (anonymous):

So 6x+2a=0? But how do you solve for a?

jimthompson5910 (jim_thompson5910):

x = 1 in this case

OpenStudy (anonymous):

Because of (1, -6)?

jimthompson5910 (jim_thompson5910):

yep, that's the point of inflection

OpenStudy (anonymous):

Then what do you do?

jimthompson5910 (jim_thompson5910):

6x+2a=0 6(1)+2a=0 6 + 2a = 0 solve for a

jimthompson5910 (jim_thompson5910):

Then once you know the value of 'a', you plug it into y=x^3+ax^2+bx-4

jimthompson5910 (jim_thompson5910):

Then you plug (1,-6) into y=x^3+ax^2+bx-4 and solve for b

jimthompson5910 (jim_thompson5910):

This works because (1,-6) is a point on y=x^3+ax^2+bx-4

OpenStudy (anonymous):

Let me try. Don't go away.

jimthompson5910 (jim_thompson5910):

alright

OpenStudy (anonymous):

Okay, I got the right answer, which is 0. Thanks.

jimthompson5910 (jim_thompson5910):

you're welcome